Ticket Number: HEA-NOC/20001018-02 Ticket Status: UPDATE
Ticket Type: Unscheduled Ticket Source: HEA-NOC
Ticket Scope: Line Site/Line: HEAnet <==> NCIR
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 20001018 13:14:00 UTC+1
Ticket Closed: Problem End:
13:14:00 NCI register as down on monitoring tools.
Traffic to and from NCI.
Mail for ncirl.ie is backed up by Kerna.
Reported problem to Eircom.
Ticket #1513 issued.
15.15 Contact Eircom for an update. They report that they will have to send an engineer out on site.
Thursday 19-10-2000 9:40 UTC. Rang Eircom for an update they say there is a break in the cable near NCI in the Sandyford Road end and it could take an hour to 24 hours to repair depending on the fault.
Time to Fix:
Left a message on Martin Moriarty's voice mail